What are the wavelengths of television station which transmits vision on 500 MHz sound
on 505 MHz respectively? (Speed of light= c= 3×108 m/s). Find (a) wavelength at which television
transmits vision and (b) wavelength at which television transmit sound.
1
Expert's answer
2020-03-13T11:03:40-0400

"c=v\lambda\\\lambda=\dfrac{c}{v}"


(a) wavelength at which television transmits vision, "\lambda=\dfrac{c}{v}=\dfrac{3\times10^8}{500\times10^6}=0.6 m"


(b) wavelength at which television transmits sound, "\lambda=\dfrac{c}{v}=\dfrac{3\times10^8}{505\times10^6}=0.59m"